OpenStudy (anonymous):

hich shows the expression below simplified? (6 × 10-7) × 0.006 3.6 × 10-10 1.2 × 1021 3.6 × 1021 3.6 × 10-9

OpenStudy (anonymous):

\[0.006=6\times 10^{-3}\]

OpenStudy (anonymous):

\[6\times 10^{-7}\times 6\times 10^{-3}=36\times 10^{-3-7}=36\times 10^{-10}\] which in scientific notation is \[3.6\times 10^{-9}\]

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

yw